OpenStudy (nathanjhw):

Let C(t) be the number of cougars on an island at time t years (where t > 0). The number of cougars is increasing at a rate directly proportional to 3500 . C(t). Also, C(0) = 1000, and C(5) = 2000.

OpenStudy (nathanjhw):

A.Find the limit as t tends to infinity of C(t) , and explain its meaning. B.On the axes provided, draw a graph showing the number of cougars as a function of time:

OpenStudy (help_people):

@perl may you help me?

OpenStudy (perl):

I think you have a typo in your question. The number of cougars is increasing at a rate directly proportional to `3500 - C(t).` The directions can be translated into the following differential equation. $$ \Large \rm { \frac{dC}{dt} = k \cdot (3500 -C(t)) }$$

OpenStudy (help_people):

@perl may you help me?

OpenStudy (perl):

we can solve that differential equation

OpenStudy (perl):

it is seperable

OpenStudy (nathanjhw):

How do we find the limit from there?

OpenStudy (perl):

$$ \Large \rm { \frac{dC}{dt} = k \cdot (3500 -C(t))\\~\\ dC = k \cdot (3500 -C(t))\cdot dt \\ ~\\ \frac{dC}{3500 -C(t)} = k \cdot dt \\ ~\\ \int \frac{dC}{3500 -C(t)} = \int k \cdot dt \\ ~\\ -\ln | 3500 - C(t) | = k\cdot t + h \\ ~ \\ solve~ for~constants~ k~ and~ h } $$

OpenStudy (perl):

does that make sense up to here?

OpenStudy (nathanjhw):

I understand how to get there, but I'm a little confused where to go from there.

OpenStudy (perl):

now we plug in our initial conditions

OpenStudy (nathanjhw):

Is that 1000 and 2000?

OpenStudy (nathanjhw):

@perl

OpenStudy (perl):

right plug those in

OpenStudy (nathanjhw):

I got -log(2500) for h

OpenStudy (nathanjhw):

k = 1.02

OpenStudy (nathanjhw):

h= -7.82 h= 1.02

OpenStudy (perl):

this is an easier expression to plug in ,

OpenStudy (nathanjhw):

Well are my solutions still correct?

OpenStudy (perl):

$$ \Large \rm { \frac{dC}{dt} = k \cdot (3500 -C(t))\\~\\ dC = k \cdot (3500 -C(t))\cdot dt \\ ~\\ \frac{dC}{3500 -C(t)} = k \cdot dt \\ ~\\ \int \frac{dC}{3500 -C(t)} = \int k \cdot dt \\ ~\\ -\ln | 3500 - C(t) | = k\cdot t + h\\ \ln | 3500 - C(t) | = -k\cdot t + h'\\~\\ e^{\ln | 3500 - C(t) |} =e^{ -k\cdot t + h'}\\ \cancel e^{\cancel \ln | 3500 - C(t) |} =e^{ -k\cdot t + h'}\\ 3500 - C(t) =e^{ -k\cdot t + h'\\}\\ 3500 - C(t) =e^{ -k\cdot t} \cdot e^{ h'}\\ 3500 - C(t) =Ae^{ -k\cdot t} \\ C(t)= 3500 - Ae^{ -k\cdot t} \\~\\ 1000 = C(0)= 3500 - Ae^{ -k\cdot 0} \\ \\ \therefore \\ 1000 = 3500 - A\cdot 1 \\ A = 3500 - 1000 =2500 } $$

OpenStudy (perl):

yes they are correct, but you are introducing decimal error , which i want to avoid

OpenStudy (nathanjhw):

Should I exclude decimals or go farther past the decimal?

OpenStudy (nathanjhw):

I guess it doesn't matter, just keep going with what you were doing.

OpenStudy (nathanjhw):

Is that the limit?

OpenStudy (perl):

that is k

OpenStudy (nathanjhw):

So now we need to find h?

OpenStudy (perl):

$$ \large \rm { \frac{dC}{dt} = k \cdot (3500 -C(t))\\~\\ dC = k \cdot (3500 -C(t))\cdot dt \\ ~\\ \frac{dC}{3500 -C(t)} = k \cdot dt \\ ~\\ \int \frac{dC}{3500 -C(t)} = \int k \cdot dt \\ ~\\ -\ln | 3500 - C(t) | = k\cdot t + h\\ \ln | 3500 - C(t) | = -k\cdot t + h'\\~\\ e^{\ln | 3500 - C(t) |} =e^{ -k\cdot t + h'}\\ \cancel e^{\cancel \ln | 3500 - C(t) |} =e^{ -k\cdot t + h'}\\ 3500 - C(t) =e^{ -k\cdot t + h'\\}\\ 3500 - C(t) =e^{ -k\cdot t} \cdot e^{ h'}\\ 3500 - C(t) =Ae^{ -k\cdot t} \\ C(t)= 3500 - Ae^{ -k\cdot t} \\~\\ 1000 = C(0)= 3500 - Ae^{ -k\cdot 0} \\ \\ \therefore \\ 1000 = 3500 - A\cdot 1 \\ A = 3500 - 1000 =2500 \\~\\ C(t)= 3500 - 2500e^{ -k\cdot t} \\~\\ 2000 = C(5) = 3500 - 2500e^{ -k\cdot 5} \\~\\ \\ \therefore \\~\\ 2000 = 3500 - 2500e^{ -k\cdot 5} \\~\\ \frac{2000-3500}{-2500} = e^{ -k\cdot 5} \\~\\ \frac 3 5 = e^{ -k\cdot 5} \\~\\ \ln (\frac 3 5) = \ln (e^{ -k\cdot 5}) \\~\\ \ln (\frac 3 5) = -k\cdot 5\\~\\ \frac{\ln ( \frac {3}{5}) }{-5} = k\\~\\ k \approx 0.102165\\~\\ C(t)= 3500 - 2500e^{ -.1022 t} \\~\\ } $$

OpenStudy (perl):

the limit you plug in infinity, and you should get

OpenStudy (nathanjhw):

t = -3.29229

OpenStudy (perl):

$$ \large \rm { \frac{dC}{dt} = k \cdot (3500 -C(t))\\~\\ dC = k \cdot (3500 -C(t))\cdot dt \\ ~\\ \frac{dC}{3500 -C(t)} = k \cdot dt \\ ~\\ \int \frac{dC}{3500 -C(t)} = \int k \cdot dt \\ ~\\ -\ln | 3500 - C(t) | = k\cdot t + h\\ \ln | 3500 - C(t) | = -k\cdot t + h'\\~\\ e^{\ln | 3500 - C(t) |} =e^{ -k\cdot t + h'}\\ \cancel e^{\cancel \ln | 3500 - C(t) |} =e^{ -k\cdot t + h'}\\ 3500 - C(t) =e^{ -k\cdot t + h'\\}\\ 3500 - C(t) =e^{ -k\cdot t} \cdot e^{ h'}\\ 3500 - C(t) =Ae^{ -k\cdot t} \\ C(t)= 3500 - Ae^{ -k\cdot t} \\~\\ 1000 = C(0)= 3500 - Ae^{ -k\cdot 0} \\ \\ \therefore \\ 1000 = 3500 - A\cdot 1 \\ A = 3500 - 1000 =2500 \\~\\ C(t)= 3500 - 2500e^{ -k\cdot t} \\~\\ 2000 = C(5) = 3500 - 2500e^{ -k\cdot 5} \\~\\ \\ \therefore \\~\\ 2000 = 3500 - 2500e^{ -k\cdot 5} \\~\\ \frac{2000-3500}{-2500} = e^{ -k\cdot 5} \\~\\ \frac 3 5 = e^{ -k\cdot 5} \\~\\ \ln (\frac 3 5) = \ln (e^{ -k\cdot 5}) \\~\\ \ln (\frac 3 5) = -k\cdot 5\\~\\ \frac{\ln ( \frac {3}{5}) }{-5} = k\\~\\ k \approx .102165 C(t)= 3500 - 2500e^{ -.1022\cdot t} \\~\\ \lim_{t \to \infty} C(t) = 3500 - 2500e^{ -.1022\cdot \infty} \\~\\ } $$

OpenStudy (perl):

what is the limit of e^{-infinity} ?

OpenStudy (nathanjhw):

0

OpenStudy (perl):

right

OpenStudy (perl):

3500 cougars is the maximum capacity or maximum population this island can reach

OpenStudy (nathanjhw):

Wow, that was harder than I thought it was going to be.

OpenStudy (nathanjhw):

I'm pretty sure I understand all that though.

OpenStudy (perl):

$$ \large \rm { \frac{dC}{dt} = k \cdot (3500 -C(t))\\~\\ dC = k \cdot (3500 -C(t))\cdot dt \\ ~\\ \frac{dC}{3500 -C(t)} = k \cdot dt \\ ~\\ \int \frac{dC}{3500 -C(t)} = \int k \cdot dt \\ ~\\ -\ln | 3500 - C(t) | = k\cdot t + h\\ \ln | 3500 - C(t) | = -k\cdot t + h'\\~\\ e^{\ln | 3500 - C(t) |} =e^{ -k\cdot t + h'}\\ \cancel e^{\cancel \ln | 3500 - C(t) |} =e^{ -k\cdot t + h'}\\ 3500 - C(t) =e^{ -k\cdot t + h'\\}\\ 3500 - C(t) =e^{ -k\cdot t} \cdot e^{ h'}\\ 3500 - C(t) =Ae^{ -k\cdot t} \\ C(t)= 3500 - Ae^{ -k\cdot t} \\~\\ 1000 = C(0)= 3500 - Ae^{ -k\cdot 0} \\ \\ \therefore \\ 1000 = 3500 - A\cdot 1 \\ A = 3500 - 1000 =2500 \\~\\ C(t)= 3500 - 2500e^{ -k\cdot t} \\~\\ 2000 = C(5) = 3500 - 2500e^{ -k\cdot 5} \\~\\ \\ \therefore \\~\\ 2000 = 3500 - 2500e^{ -k\cdot 5} \\~\\ \frac{2000-3500}{-2500} = e^{ -k\cdot 5} \\~\\ \frac 3 5 = e^{ -k\cdot 5} \\~\\ \ln (\frac 3 5) = \ln (e^{ -k\cdot 5}) \\~\\ \ln (\frac 3 5) = -k\cdot 5\\~\\ \frac{\ln ( \frac {3}{5}) }{-5} = k\\~\\ k \approx .102165\\~\\ C(t)= 3500 - 2500e^{ -.1022\cdot t} \\~\\ \lim_{t \to \infty} C(t) = 3500 - 2500e^{ -.1022\cdot \infty} \\ = 3500 - 2500\cdot 0 = 3500 } $$

OpenStudy (nathanjhw):

I'll be sure to look over all that work.

OpenStudy (nathanjhw):

However, can you help me with B right now?

OpenStudy (perl):



OpenStudy (perl):

the dashed line is not part of the curve, that is an asymptote , as \(\bf t \to \infty \)

OpenStudy (nathanjhw):

Yeah, I understand that.

OpenStudy (nathanjhw):

You've been a real help. Thanks so much!